A company makes unicycles,bicycles and tricycles. Last week they made 88 more bicycles than unicycles, and 5 times as many tricycles as unicycles. If they made 40 more bicycles than tricycles, how many unicycles did they make?
Found 2 solutions by sarah_adam, priya:
Answer by sarah_adam(201)

Let Number of Bicycles,Unicycles and Tricycles be B,U and T respectively
given :
B = 88+ U
T = 5U
B = 40+T
88+U = 40 +T
88+U = 40 + 5U
5U - U = 88 - 40
4U = 44
U = 12
Number of unicycles are 12

Answer by priya(51)

First try to form equations, with the informations given in the question.
let U be unicycles,B be bicycles and T be tricycles.
then,
eq 1. B = 88 + U
eq 2. T = 5 * U

eq 3. B = 40 + T (GIVEN )

compare eq 1 & eq 2,we get
88 + U = 40 + T
NOW FROM eq 2. , substitute T = 5*U, we get
88 + U = 40 + (5 * U)
88 - 40 = 5U - U
48 = 4U
Therefore U = 12.SO THE NUMBER OF UNICYCLES IS 12.
WANT TO LEARN ONLINE..VISIT MY SITE!!
PRIYA